4x+2y=-18
5x+y=-27
What is the solution to the system of equations using linear combination or substitution.

Answer by mananth(16946)

4 x + 2 y = -18 .............1
5 x + 1 y = -27 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by -2
4 x 2 y = -18
-10 x -2 y = 54
Add the two equations
-6 x = 36
/ -6
x = -6
plug value of x in (1)
4 x + 2 y = -18
-24 + 2 y = -18
2 y = -18 + 24
2 y = 6
y = 3
solution (-6,3)